# The Standard Model lifetime is slightly shorter

### v1

@@ Abstract
We revisit the decay rate of the electroweak vacuum in the Standard Model with the full one-loop prefactor. We focus on the gauge degrees of freedom and derive the degeneracy factors appearing in the functional determinant using group theoretical arguments. Our treatment shows that the transverse modes were previously overcounted, so we revise the calculation of that part of the prefactor. The new result modifies the gauge fields’ contribution by $7\%$ and slightly decreases the previously predicted lifetime of the electroweak vacuum, which remains much longer than the age of the universe. Our discussion of the transverse mode degeneracy applies to any calculation of functional determinants involving gauge fields in four dimensions.
@@
Figure 1: Contours of $\gamma$ as a function of the Higgs $m_{h}$ and the top mass $m_{t}$. The black dot-dashed contours correspond to $\gamma=1,10^{-100},10^{-300}$ and $10^{-1000}\text{ Gyr}^{-1}\text{Gpc}^{-3}$ for $\alpha_{s}(m_{Z})=0.1179$. In the blue region $\gamma$ becomes larger than $H_{0}^{4}$, with $H_{0}$ the current Hubble parameter, while in the yellow region the EW vacuum is stable. The boundaries of these regions are given by the plain lines for $\alpha_{s}(m_{Z})=0.1179$, the dotted lines for $\alpha_{s}(m_{Z})=0.1170$, the dashed line for $\alpha_{s}(m_{Z})=0.1188$. The green region, for which $\gamma=10^{-781}\text{ Gyr}^{-1}\text{Gpc}^{-3}$ at the center, shows the experimentally measured values of the Higgs and top masses with their $1\sigma$ (inside), $2\sigma$ (middle) and $3\sigma$ (outside) uncertainties in quadrature.
@@
In this Letter we revisit the calculation of the gauge prefactor. We find that the transverse mode degeneracy was not properly taken into account. Once corrected, the central value of the SM rate $\gamma$ increases only slightly, from the previous $10^{-785}$ to $10^{-781}$ Gyr-1 Gpc-3. The SM vacuum lifetime remains longer than the current age of the universe and there is no occasion for anxiety Coleman (1977).
@@
Vector field basis. For the spin-1 field, the total angular momentum is $J_{\mu\nu}=L_{\mu\nu}+S_{\mu\nu}$, where $[S_{\mu\nu}]_{\rho\sigma}=-i(\delta_{\mu\rho}\delta_{\nu\sigma}-\delta_{\mu\sigma}\delta_{\nu\rho})$ are the generators acting on the spin-component. The vector field transforms like $A_{\mu}(x)\to R_{\mu\nu}\,A_{\nu}(R^{-1}x)$, where $R_{\mu\nu}$ is the vector rotation matrix. The representation of $A_{\mu}(x)$ is isomorphic to the tensor product of a spin-1 component $(1/2,1/2)$, with an orbital component, transforming as $\oplus_{l=0}^{\infty}(l/2,l/2)$. To understand the $SO(4)$ irreps $(j_{A},j_{B})$ of the vector field, we expand the tensor product:
@@
To go from (5) to (6) we simply re-labeled and rearranged the sum. In (6) we included a subscript for each $(j_{A},j_{B})$ multiplet to track the orbital angular momentum quantum number $l$. We call the multiplets with $j_{A}=j_{B}$ the diagonal modes ${\cal D}^{\pm}_{j}$, corresponding to $(j/2,j/2)_{j\pm 1}$ (note ${\cal D}^{-}_{0}$ is zero); they have eigenvalue $j(j+2)$ under $J^{2}$, and degeneracy $d_{j}^{\cal D}\equiv{\rm dim}\left(\frac{j}{2},\frac{j}{2}\right)=\left(j+1\right)^{2}$. These are the same total angular momentum quantum numbers as the NGB scalars, hence the fluctuation operator matrix mixes these states. We identify the multiplets having $j_{A}=j_{B}\pm 1$ with the transverse modes Higuchi (1991), and call them ${\cal T}^{\pm}_{j}$ for $((j\pm 1)/2,(j\mp 1)/2)_{j}$. Their $J^{2}$ eigenvalue is $(j+1)^{2}$, while the degeneracy factor is
@@
where the real field $A_{\mu}$ is expanded with complex basis, without affecting the degeneracy, as follows
@@
This choice has the virtue of containing familiar objects that are in direct correspondence with the group theory construction of FIG. 2. The $(\tilde{\sigma}_{\mu})_{m_{s_{A}}m_{s_{B}}}$ matrices correspond to the $(1/2,1/2)$ object that ensures the proper transformation under rotations, and are given by $\tilde{\sigma}_{\mu}=\varepsilon\cdot\sigma_{\mu}$, where $\varepsilon$ is the two-dimensional Levi-Civita symbol, $\sigma_{1,2,3}$ are the usual Pauli matrices, and $\sigma_{4}=\text{diag}(i,i)$. The entries of $\tilde{\sigma}$ are ordered such that $m_{s}=1/2$ comes before $m_{s}=-1/2$. $C^{j_{A},m_{A}}_{l_{A}m_{l_{A}}s_{A}m_{s_{A}}}$ denote the $SU(2)_{A}$ Clebsch-Gordan coefficients (and the same for $B$). These are non-zero only when $|2j_{A,B}-l|=1$, $m_{A}=m_{l_{A}}+m_{s_{A}}$ and $m_{B}=m_{l_{B}}+m_{s_{B}}$. We shall suppress the $m$ indices from here on for brevity; they label the states with the same fluctuation operator and their summation only appears as the degeneracy factor. Then, the non-zero components of $V^{j_{A}j_{B}}_{l;\mu}$ are $V^{\frac{j}{2}\frac{j}{2}}_{l=j\pm 1;\mu}$ and $V^{\frac{j\pm 1}{2}\frac{j\mp 1}{2}}_{l=j;\mu}$, corresponding to the basis of ${\cal D}_{j}^{\pm}$ and ${\cal T}_{j}^{\pm}$. These components are eigenfunctions of $J^{2}$, with the eigenvalues quoted earlier for ${\cal D}_{j}^{\pm}$ and ${\cal T}_{j}^{\pm}$. Each component is an eigenfunction of $L^{2}$ that acts only on $Y_{lm_{l_{A}}m_{l_{B}}}$.
@@
With this decomposition, the prefactor in (2) factorizes into ${\cal A}^{(A,\varphi)}={\cal A}^{({\cal D},\varphi)}{\cal A}^{({\cal T})}$. The first term, ${\cal A}^{({\cal D},\varphi)}$, was already computed correctly in the previous literature Isidori et al. (2001); Andreassen et al. (2018); Chigusa et al. (2017). Therefore we only deal with the second one:
@@
  • In previous calculations Isidori et al. (2001); Andreassen et al. (2018); Chigusa et al. (2017), the degeneracy factor for the transverse modes, ${\cal T}_{j}^{\pm}$, was erroneously equal to $(j+1)^{2}$. The correct one, as we have shown in (7), is $d_{j}^{\cal T}=j(j+2)$. This leads to a slightly different result for the prefactor ${\cal A}^{(\cal T)}$. We will shortly revise that calculation.
  • For the diagonal modes ${\cal D}^{\pm}_{j}$, our basis functions have one-to-one correspondence to the ones given in Isidori et al. (2001), which are then used in Andreassen et al. (2018); Chigusa et al. (2017, 2018). We differ for the transverse modes ${\cal T}^{\pm}_{j}$. In the supplemental material we show that their basis functions are linearly dependent and do not span the entire space. Despite the degeneracy factor and completeness issues, the operator $S^{\prime\prime({\cal T})}_{j}$ in Isidori et al. (2001); Andreassen et al. (2018); Chigusa et al. (2017) is correct.
@@
We use the current values of the SM parameters and their $1\sigma$ errors from Workman et al. (2022): $m_{h}=125.25\pm 0.17\text{ GeV}$, $m_{t}=172.69\pm 0.30\text{ GeV}$ and $\alpha_{s}(m_{Z})=0.1179\pm 0.0009$. Our final decay rate is
$\log_{10}\frac{\gamma}{\rm Gyr\,Gpc^{3}}=-781^{+45+154+180}_{-49-220-274}\,,$ (17)
where the first, second, and third errors are calculated by changing the Higgs mass, the top mass, and the strong coupling by $1\sigma$, respectively. The variations are summarized in FIG. 1, which is similar to the one in the Addendum of Chigusa et al. (2017), where they used the same SM inputs.
Comparing to the most recent previous result Chigusa et al. (2017), we find $\gamma/\gamma_{\rm prev}\approx 10^{4}$, using the central values of the SM parameters. The change of the central value is much smaller than the uncertainties in (17). However, it is still larger than the theoretical uncertainty ($\approx 10^{3}$), evaluated by setting the renormalization scale to $\mu=2/R$ and $\mu=1/(2R)$ instead of $\mu=1/R$.
@@
Appendix A Previous vector field basis
@@
where $A^{(B,L,T1,T2)}_{jm_{A}m_{B}}(\rho)$ are the radial parts, functions of the Euclidean radius $\rho$ only. Here, the labels $j,m_{A}$ and $m_{B}$ correspond to the labels $j,m$ and $m^{\prime}$ introduced in Isidori et al. (2001), which indicate the hyperspherical harmonics that the basis functions are constructed from. The orbital parts do not depend on $\rho$ and are only functions of $\hat{x}$, or the three angles in 4D. We separate them into $B$ modes, proportional to the ‘breathing’ direction $\hat{x}_{\mu}$, and longitudinal-like $L$ modes along the momentum $\rho\partial_{\mu}$:
@@
In particular, the transverse modes $T^{1}$ and $T^{2}$ only cover a space with dimension $2j(j+2)-j$ for any $j>0$. This is neither equal to the naïve $2(j+1)^{2}$, nor to the correct $2j(j+2)$, which we will obtain with group theoretical arguments. Thus, the basis functions in (18) have to be dependent and incomplete. Let us start with the $B$ and $L$ modes, which are orthogonal to each other. The $j=0$ mode is constructed from $Y_{000}(\hat{x})=1/\sqrt{4\pi}$. Therefore, only $B_{\mu}^{000}$ is non zero, while $L_{\mu}^{000}$ and $T^{i;000}_{\mu}$ vanish because they contain a derivative hitting the constant $Y_{000}$. Also, $B_{\mu}^{000}$ is proportional to $V^{00}_{l=1;\mu}$ [see (9)], which corresponds to the singlet state with $j=0$. Next, consider $j\geq 1$. Both $B^{jm_{A}m_{B}}_{\mu}$ and $L^{jm_{A}m_{B}}_{\mu}$ are non-zero for all $m_{A}$ and $m_{B}$. Note that the $\hat{x}^{\mu}$ and $\partial_{\mu}$ operators commute with $J_{\mu\nu}$. This implies that $B^{jm_{A}m_{B}}_{\mu}$ and $L^{jm_{A}m_{B}}_{\mu}$ have the total angular momentum quantum numbers $j_{A}=j_{B}=j/2$, and are orthogonal to each other. They are eigenfunctions of $J^{2}$ with eigenvalue $j(j+2)$, and have the same degeneracy factor as the hyperspherical harmonics, $(j+1)^{2}$.
@@
We have recovered the result of the group theoretical argument in the main text. It follows that the modes $B^{jm_{A}m_{B}}_{\mu}$ and $L^{jm_{A}m_{B}}_{\mu}$ cover exactly the same space spanned by our ${\cal D}_{j}^{\pm}$ basis, which is a $2(j+1)^{2}$ dimensional space. Unlike $V^{j_{A}j_{B}}_{l;\mu}$, the $B^{jm_{A}m_{B}}_{\mu}$ and $L^{jm_{A}m_{B}}_{\mu}$ are still not eigenstates of $L^{2}$ and it is diagonalized by taking linear combinations,
@@
Appendix B Transverse gauge modes
First note that the $T^{i;jm_{A}m_{B}}_{\mu}$ are not eigenstates of $J_{A3,B3}$, the third component of the $SU(2)_{A}$ and $SU(2)_{B}$ total angular momentum. Hence, the labels $m_{A}$ and $m_{B}$ do not correspond to the eigenvalues of $J_{A3}$ and $J_{B3}$, they are just inherited from the labels of spherical harmonics. Thus having different $m_{A}$ and $m_{B}$ does not ensure that the functions are independent. The $J^{2}$ eigenvalue for $T^{i;jm_{A}m_{B}}_{\mu}$ is given by $(j+1)^{2}$, therefore they have to be related to the ${\cal T}^{\pm}_{j}$ modes. However, we know from the discussion in the main text that there exist only $2j(j+2)$ independent basis functions for ${\cal T}^{\pm}_{j}$. Let us then analyze the space spanned by $T^{1,2;jm_{A}m_{B}}_{\mu}$.
Computing the degeneracy of $T^{1}$. We invoke again the rank of the Gram matrix. First, though, we need some preliminaries. Without loss of generality we take $V^{(1)}_{\mu}=\delta_{\mu 4}$ and $V^{(2)}_{\mu}=\delta_{\mu 3}$ as the two independent vectors in (20). Then we define the $SU(2)_{A}$ and $SU(2)_{B}$ orbital angular momentum operators as

### v2

# Revising the full one-loop gauge prefactor in electroweak vacuum stability

@@ Abstract
We revisit the decay rate of the electroweak vacuum in the Standard Model with the full one-loop prefactor. We focus on the gauge degrees of freedom and derive the degeneracy factors appearing in the functional determinant using group theoretical arguments. Our treatment shows that the transverse modes were previously overcounted, so we revise the calculation of that part of the prefactor. The new result modifies the gauge fields’ contribution by $6\%$ and slightly decreases the previously predicted lifetime of the electroweak vacuum, which remains much longer than the age of the universe. Our discussion of the transverse mode degeneracy applies to any calculation of functional determinants involving gauge fields in four dimensions.
@@
Figure 1: Contours of $\gamma$ as a function of $\alpha_{s}(m_{Z})$ and the top mass $m_{t}$. The black dot-dashed contours correspond to $\gamma=1,10^{-100},10^{-300}$ and $10^{-1000}\text{ Gyr}^{-1}\text{Gpc}^{-3}$ for the Higgs mass $m_{h}=125.20$ GeV. In the blue region $\gamma$ becomes larger than $H_{0}^{4}$, with $H_{0}$ the current Hubble parameter, while in the yellow region the EW vacuum is stable. The boundaries of these regions are given by the plain lines for $m_{h}=125.20$ GeV, the dotted lines for $m_{h}=125.09$ GeV, the dashed line for $m_{h}=125.31$ GeV. The green region, for which $\gamma=10^{-871}\text{ Gyr}^{-1}\text{Gpc}^{-3}$ at the center, shows the experimentally measured values of $\alpha_{s}(m_{Z})$ and the top mass with their $1\sigma$ (inside), $2\sigma$ (middle) and $3\sigma$ (outside) uncertainties in quadrature.
@@
In this Letter we revisit the calculation of the gauge prefactor. We find that the transverse mode degeneracy was not properly taken into account. Once corrected, the central value of the SM rate $\gamma$ increases only slightly, from $10^{-877}$ to $10^{-871}$ Gyr-1 Gpc-3. The SM vacuum lifetime remains longer than the current age of the universe and there is no occasion for anxiety Coleman (1977).
@@
Vector field basis. For the spin-1 field, the total angular momentum is $J_{\mu\nu}=L_{\mu\nu}+S_{\mu\nu}$, where $[S_{\mu\nu}]_{\rho\sigma}=-i(\delta_{\mu\rho}\delta_{\nu\sigma}-\delta_{\mu\sigma}\delta_{\nu\rho})$ are the generators acting on the spin-component. Equivalently, the vector field transforms like $A_{\mu}(x)\to R_{\mu\nu}\,A_{\nu}(R^{-1}x)$, where $R_{\mu\nu}$ is the vector rotation matrix. The representation of $A_{\mu}(x)$ is isomorphic to the tensor product of a spin-1 component $(1/2,1/2)$, with an orbital component, transforming as $\oplus_{l=0}^{\infty}(l/2,l/2)$. To understand the $SO(4)$ irreps $(j_{A},j_{B})$ of the vector field, we expand the tensor product:
@@
To go from (5) to (6) we simply re-labeled and rearranged the sum. In (6) we included a subscript for each $(j_{A},j_{B})$ multiplet to track the orbital angular momentum quantum number $l$. We call the multiplets with $j_{A}=j_{B}$ the diagonal modes ${\cal D}^{\pm}_{j}$, corresponding to $(j/2,j/2)_{j\pm 1}$ (note ${\cal D}^{-}_{0}$ is zero); they have eigenvalue $j(j+2)$ under $J^{2}$, and degeneracy $d_{j}^{\cal D}\equiv{\rm dim}\left(\frac{j}{2},\frac{j}{2}\right)=\left(j+1\right)^{2}$. These are the same total angular momentum quantum numbers as the NGB scalars, hence in general the fluctuation operator matrix mixes these states. We identify the multiplets having $j_{A}=j_{B}\pm 1$ with the transverse modes Higuchi (1991), and call them ${\cal T}^{\pm}_{j}$ for $((j\pm 1)/2,(j\mp 1)/2)_{j}$. Their $J^{2}$ eigenvalue is $(j+1)^{2}$, while the degeneracy factor is
@@
where the real field $A_{\mu}$ is expanded with a complex basis, without affecting the degeneracy, as follows
@@
This choice has the virtue of containing familiar objects that are in direct correspondence with the group theory construction of FIG. 2. The $(\tilde{\sigma}_{\mu})_{m_{s_{A}}m_{s_{B}}}$ matrices correspond to the $(1/2,1/2)$ object that ensures the proper transformation under rotations, and are given by $\tilde{\sigma}_{\mu}=\varepsilon\cdot\sigma_{\mu}$, where $\varepsilon$ is the two-dimensional Levi-Civita symbol, $\sigma_{1,2,3}$ are the usual Pauli matrices, and $\sigma_{4}=\text{diag}(i,i)$. The entries of $\tilde{\sigma}$ are ordered such that $m_{s}=1/2$ comes before $m_{s}=-1/2$. $C^{j_{A},m_{A}}_{l_{A}m_{l_{A}}s_{A}m_{s_{A}}}$ denote the $SU(2)_{A}$ Clebsch-Gordan coefficients (and the same for $B$). These are non-zero only when $|2j_{A,B}-l|=1$, $m_{A}=m_{l_{A}}+m_{s_{A}}$ and $m_{B}=m_{l_{B}}+m_{s_{B}}$. We shall suppress the $m$ indices from here on for brevity; they label states with the same fluctuation operator and their summation only appears as the degeneracy factor. Then, the non-zero components of $V^{j_{A}j_{B}}_{l;\mu}$ are $V^{\frac{j}{2}\frac{j}{2}}_{l=j\pm 1;\mu}$ and $V^{\frac{j\pm 1}{2}\frac{j\mp 1}{2}}_{l=j;\mu}$, corresponding to the basis of ${\cal D}_{j}^{\pm}$ and ${\cal T}_{j}^{\pm}$. These components are eigenfunctions of $J^{2}$, with the eigenvalues quoted earlier for ${\cal D}_{j}^{\pm}$ and ${\cal T}_{j}^{\pm}$. Each component is an eigenfunction of $L^{2}$ that acts only on $Y_{lm_{l_{A}}m_{l_{B}}}$.
@@
With this decomposition, the prefactor in (2) factorizes into ${\cal A}^{(A,\varphi)}={\cal A}^{({\cal D},\varphi)}{\cal A}^{({\cal T})}$. The first term ${\cal A}^{({\cal D},\varphi)}$ was already computed correctly in the previous literature Isidori et al. (2001); Andreassen et al. (2018); Chigusa et al. (2017). Therefore we only deal with the second one:
@@
  • In previous calculations Isidori et al. (2001); Andreassen et al. (2018); Chigusa et al. (2017), the degeneracy factor for the transverse modes ${\cal T}_{j}^{\pm}$ was erroneously set equal to $(j+1)^{2}$. The correct one, as we have shown in (7), is $d_{j}^{\cal T}=j(j+2)$. This leads to a slightly different result for the prefactor ${\cal A}^{(\cal T)}$. We will shortly revise that calculation.
  • For the diagonal modes ${\cal D}^{\pm}_{j}$, our basis functions have one-to-one correspondence to the ones given in Isidori et al. (2001), which are then used in Andreassen et al. (2018); Chigusa et al. (2017, 2018). We differ for the transverse modes ${\cal T}^{\pm}_{j}$. In the appendix of the end matter we give an explicit demonstration of why the gauge boson basis in Isidori et al. (2001) has inconsistencies. In the supplemental material we furhter elaborate and show that their basis functions are linearly dependent and do not span the entire space. Despite the degeneracy factor and completeness issues, the operator $S^{\prime\prime({\cal T})}_{j}$ in Isidori et al. (2001); Andreassen et al. (2018); Chigusa et al. (2017) is correct.
@@
We use the current values of the SM parameters and their $1\sigma$ errors from Navas et al. (2024): $m_{h}=125.20\pm 0.11\text{ GeV}$, $m_{t}=172.57\pm 0.29\text{ GeV}$ and $\alpha_{s}(m_{Z})=0.1180\pm 0.0009$. Our final decay rate is
$\log_{10}\frac{\gamma}{\rm Gyr^{-1}\,Gpc^{-3}}=-871^{+35+175+209}_{-37-253-330}\,,$ (17)
where the first, second, and third errors are calculated by changing the Higgs mass, the top mass, and the strong coupling by $1\sigma$, respectively. The variations are summarized in FIG. 1.
Comparing to the most recent previous result Chigusa et al. (2017), we find $\gamma/\gamma_{\rm prev}\approx 10^{6}$, using the central values of the SM parameters. The change of the central value is much smaller than the uncertainties in (17). However, it is still larger than the theoretical uncertainty ($\approx 10^{3}$), evaluated by setting the renormalization scale to $\mu=2/R$ and $\mu=1/(2R)$ instead of $\mu=1/R$.
@@
I Previous vector field basis
@@
where $A^{(B,L,T1,T2)}_{jm_{A}m_{B}}(\rho)$ are the radial parts, functions of the Euclidean radius $\rho$ only. Here, the labels $j,m_{A}$ and $m_{B}$ are exactly the labels $j,m$ and $m^{\prime}$ introduced in Isidori et al. (2001), which indicate the hyperspherical harmonics that the basis functions are constructed from. The orbital parts do not depend on $\rho$ and are only functions of $\hat{x}$, or the three angles in 4D. We separate them into $B$ modes, proportional to the ‘breathing’ direction $\hat{x}_{\mu}$, and longitudinal-like $L$ modes along the momentum $\rho\partial_{\mu}$:
@@
In particular, the transverse modes $T^{1;j}$ and $T^{2;j}$ only cover a space with dimension $2j(j+2)-j$ for any $j>0$, which is definitely not equal to the naïve $2(j+1)^{2}$. As we will show later, the correct degeneracy is $2j(j+2)$, so the basis of Isidori et al. (2001) misses $j$ elements, meaning that it is both dependent and incomplete. Let us start with the $B$ and $L$ modes, which are orthogonal to each other. The $j=0$ mode is constructed from $Y_{000}(\hat{x})=1/\sqrt{4\pi}$. Therefore, only $B_{\mu}^{000}$ is non zero, while $L_{\mu}^{000}$ and $T^{i;000}_{\mu}$ vanish because they contain a derivative hitting the constant $Y_{000}$. Also, $B_{\mu}^{000}$ is proportional to $V^{00}_{l=1;\mu}$ [see (9)], which corresponds to the singlet state. Next, consider $j\geq 1$. Both $B^{jm_{A}m_{B}}_{\mu}$ and $L^{jm_{A}m_{B}}_{\mu}$ are non-zero for all $m_{A}$ and $m_{B}$. Note that the $\hat{x}^{\mu}$ and $\partial_{\mu}$ operators commute with $J_{\mu\nu}$. This implies that $B^{jm_{A}m_{B}}_{\mu}$ and $L^{jm_{A}m_{B}}_{\mu}$ have the total angular momentum quantum numbers $j_{A}=j_{B}=j/2$, and are orthogonal to each other. They are eigenfunctions of $J^{2}$ with eigenvalue $j(j+2)$, and have the same degeneracy factor as the hyperspherical harmonics, $(j+1)^{2}$.
@@
We have recovered the result of the group theoretical argument in the main text. It follows that the modes $B^{jm_{A}m_{B}}_{\mu}$ and $L^{jm_{A}m_{B}}_{\mu}$ cover exactly the same space spanned by our ${\cal D}_{j}^{\pm}$ basis, which is a $2(j+1)^{2}$ dimensional space. Unlike $V^{j_{A}j_{B}}_{l;\mu}$, the $B^{jm_{A}m_{B}}_{\mu}$ and $L^{jm_{A}m_{B}}_{\mu}$ are still not eigenstates of $L^{2}$ and are diagonalized by taking linear combinations,
@@
II Transverse gauge modes
First note that the $T^{i;jm_{A}m_{B}}_{\mu}$ of (23) are not eigenstates of $J_{A3,B3}$, the third component of the $SU(2)_{A}$ and $SU(2)_{B}$ total angular momentum. Hence, the labels $m_{A}$ and $m_{B}$ do not correspond to the eigenvalues of $J_{A3}$ and $J_{B3}$, they are just inherited from the labels of spherical harmonics. Thus having different $m_{A}$ and $m_{B}$ does not ensure that the functions are independent. The $J^{2}$ eigenvalue for $T^{i;jm_{A}m_{B}}_{\mu}$ is given by $(j+1)^{2}$, therefore they have to be related to the ${\cal T}^{\pm}_{j}$ modes. However, we know from the discussion in the main text that there exist only $2j(j+2)$ independent basis functions for ${\cal T}^{\pm}_{j}$. Let us then analyze the space spanned by $T^{1,2;jm_{A}m_{B}}_{\mu}$.
Computing the degeneracy of $T^{1}$. We invoke again the rank of the Gram matrix. First we need some preliminaries. Without loss of generality we take $V^{(1)}_{\mu}=\delta_{\mu 4}$ and $V^{(2)}_{\mu}=\delta_{\mu 3}$ as the two independent vectors in (25). Then we define the $SU(2)_{A}$ and $SU(2)_{B}$ orbital angular momentum operators as